OpenStudy (anonymous):

What is the solution to the system of equations? x + y = 2 x - y = 4

OpenStudy (anonymous):

This one is easy because you can simply solve this vertically. y and -y cancel each other out, so you're left with 2x = 6. Divide each side by 2 and you're left with 3. x = 3 Now plug in x into one of the original equations. x + y = 2 3 + y = 2 -3 -3 y = -1 So x = 3 and y = -1 You can plug both in to check your answer. x + y = 2 3 + (-1) = 2 2 = 2 Yes. x - y = 4 3 - (-1) = 4 3 + 1 = 4 4=4 Yes

OpenStudy (anonymous):

Did that help? (:

OpenStudy (anonymous):

Yes, That helped! Thanks :)

OpenStudy (anonymous):

You're welcome.